# References Matter: Investigating the Impact of Reference Set Variation on Summarization Evaluation

### v2

@@ 1 Introduction
The assumption behind the use of reference-based metrics is that system outputs that are more similar to the reference(s) are better, due to their “human-likeness” Gehrmann et al. (2023). However, the significant variation in human-written summaries implies that evaluating system outputs against a single or limited set of references has inherent drawbacks. Previous research has extensively looked at correlation between metrics and human judgments in summarization Forde et al. (2024), further exploring the use of multiple references to improve such correlations Lin (2004); Belz and Reiter (2006); Fabbri et al. (2021); Tang et al. (2024) as well as the interpretability and efficiency aspects of such automatic metrics Liu et al. (2023b). However, a much less studied question is the extent to which automatic metrics are sensitive to the choice of human-written reference summaries, as shown in Figure 1. In other words, are these metrics stable across different plausible gold-standard references? If metric scores vary significantly with the selected reference(s), this variation calls into question the reliability of many evaluation practices in the field.
@@ 2 Related Work
Recent advances in Natural Language Generation (NLG) have significantly enhanced the development of summarization systems. However, their evaluation remains an open problem Celikyilmaz et al. (2021); Goyal et al. (2023). Summarization evaluation metrics are broadly categorized into reference-based and reference-free Lloret et al. (2018). Reference-based metrics compare system outputs to human-written reference summaries, relying on methods such as n-gram overlap Lin (2004); Papineni et al. (2002), embedding similarity Ng and Abrecht (2015); Zhao et al. (2019); Zhang et al. (2020), or model-based techniques Peyrard et al. (2017); Scialom et al. (2019); Yuan et al. (2021). In contrast, reference-free summarization metrics do not assume a gold standard Yuan et al. (2021); Vasilyev et al. (2020); Gao et al. (2020). Lastly, growing research leverages LLMs as evaluators, with or without references Liu et al. (2023a); Song et al. (2024); Li et al. (2024); Gigant et al. (2024).
@@
Meta-evaluation of summarization metrics typically focuses on the extent to which they can be used as a proxy for human evaluation. Reiter and Belz (2009) examined the validity of automatic scores for NLG tasks, while Rankel et al. (2013) focused on ROUGE and its correlation with humans. Peyrard (2019) showed that metrics with reasonable correlation on lower-quality outputs tend to diverge when output quality increases. Caglayan et al. (2020) demonstrated the idiosyncrasies of automatic evaluation metrics, noting that high correlation with human judgments is not sufficient to characterize their reliability. Fabbri et al. (2021) performed a large-scale meta-evaluation of summarization metrics, and found that most metrics have low correlation with human judgments on coherence, while relevance is weakly or moderately correlated.
While most existing research focused on correlation to human scores, Tang et al. (2024) addressed the challenge of evaluation when a limited number of references is available. They proposed leveraging LLMs to diversify the references, expanding the evaluation coverage and improving the correlation with humans. Their results show that increasing the number of references significantly enhances the reliability of existing evaluation metrics in terms of correlation. However, since LLM outputs tend to show less variability and follow distinct patterns compared to human-produced content Giulianelli et al. (2023); Guo et al. (2024); Shur-Ofry et al. (2024); Reinhart et al. (2025), relying on them to replace human references might introduce biases.
@@ 3 Experimental Setup
To quantify the impact of human-written references on the scores of automatic metrics, we exam multiple elements. For datasets, we use SummEval Fabbri et al. (2021), GUMSum Liu and Zeldes (2023), and DUC2004 Dang and Croft (2004), which contain multiple human-written summaries (§3.1), to assess how different reference summaries affect metric performances. Next, to assess summarization models, we use the existing outputs provided by Fabbri et al. (2021) for SummEval. As these outputs predate LLMs, we additionally collect outputs using LLMs (§3.2) for all three datasets. Lastly, to compute the correlations with humans, we use the human judgments available in SummEval and gather new human ratings for GUMSum on both human and LLM-generated summaries (§3.3). We prioritize GUMSum over DUC2004, as it includes multiple genres beyond news data. Our metric selection is outlined in §3.4. Details on data licensing and codebase are provided in Appendix A.
@@ 3.1 Human-written Summaries
Table 1 provides an overview of the three datasets.
SummEval Fabbri et al. (2021) is built on top of CNN/DM Hermann et al. (2015); Nallapati et al. (2016), containing news articles and human-written highlights. The original authors selected 100100100 instances from the test set and supplemented the existing highlights with ten additional reference summaries per instance, obtained via crowd-sourcing Kryscinski et al. (2019).
GUMSum Liu and Zeldes (2023) contains summaries created following general and genre-specific guidelines to function as a substitute for the source Nenkova and McKeown (2011). We focus on the 484848 documents in the dev and test sets, which contain five human-written summaries each Lin and Zeldes (2025) across 121212 genres.
DUC2004 Task1 Dang and Croft (2004) consists of 489489489 news documents, most with four references. The guidelines allow the summaries to be in the form of short sentences or lists of keywords. [^1] https://duc.nist.gov/duc2004/tasks.html DUC2004 references are thus extremely concise (only up to 757575 characters). The dataset has played a significant role in summarization research, being part of the annual TREC conference evaluation.
@@ 3.2 Model Outputs
Fabbri et al. (2021) collected model outputs for SummEval from 242424 extractive and abstractive summarization systems, which were SOTA between 2017 and 2019. We focus on the 161616 models for which they provided human judgments.
For all datasets, we also include summaries generated by contemporary LLMs. This is crucial given that prior studies demonstrated that evaluation metrics often show lower correlation with high-quality outputs Peyrard (2019); Alva-Manchego et al. (2021). Below, we report a similar pattern for LLMs (§4.4). For consistency purposes, we follow Lin and Zeldes (2025) and use Llama3-3B-Instruct Hermann et al. (2015), Qwen-2.5-7B-Instruct Qwen et al. (2025), Claude-3.5 Anthropic (2024), and GPT-4o OpenAI (2024). For each LLM, we generate a single summary. We emphasize LLM variety over multiple generations. Details on the generation parameters and prompts are reported in Appendix A.
@@ 3.3 Human Judgments
SummEval (Fabbri et al., 2021) contains expert judgments that assess summaries based on four criteria: coherence, consistency, fluency, and relevance, using a Likert scale of 111-555 Likert (1932).
To measure how well automatic metrics align with human judgments in different genres, and to study whether findings on pre-LLM models align with those on LLM outputs, we collect a new set of human judgments using the same criteria on the 484848 GUMSum documents. We hired three Master’s students in Computational Linguistics and tasked them to evaluate four LLM outputs (§3.2) and five human references, following Fabbri et al. (2021)’s criteria.
Summarizer Coh. Con. Flu. Rel. best worst
claude 4.750.45 4.480.65 4.820.42 4.260.71 0.170.38 0.030.18
gpt4o 4.420.61 4.610.63 4.740.45 4.070.65 0.110.32 0.120.32
Qwen 4.660.56 4.330.84 4.680.52 4.230.75 0.160.37 0.120.33
Llama3 4.710.51 4.140.97 4.780.42 4.090.86 0.120.32 0.200.40
humans 4.540.57 4.480.71 4.700.51 4.220.69 0.090.28 0.100.31
Table 2: Human judgments on GUMSum for system- versus human-written summaries.
LLM-generated and human-written summaries were anonymized and shuffled. We also asked the evaluators to pick one best and one worst summary for each document. Table 2 reports the results. Claude scored the best overall. GPT-4o gets the highest consistency but the lowest coherence and relevance, and thus the least picked LLM. Interestingly, LLM outputs typically receive higher scores than human-written references, in line with existing research Zhang et al. (2024).
@@ 3.4 Evaluation Metrics
To gain insights into recent metric usage, Figure 2 summarizes the percentage of summarization papers from recent ACL/EMNLP/INLG proceedings (detailed in Appendix C). We found that reference-based metrics are still the most popular, with ROUGE in 797979% of papers, followed by BERTScore (444444%). Our preliminary search also shows that LLM-as-a-judge evaluators are mainly (i.e., in 666666% of the cases) used without references, placing them outside the scope of this paper. Below, we describe the commonly used reference-based metrics in the surveyed papers, balancing popularity and diversity. All metrics range in 0100100100. Appendix B provides details.
@@
BLEU Papineni et al. (2002) is an n-gram metric primarily used for translation. It is precision-based and incorporates a brevity penalty. With multiple references, the n-gram count is clipped at the maximum count of n-grams in a single reference and the length of the reference closest in size to the hypothesis is considered.
@@ 4.1 Human-written Summaries are Diverse
Figure 3 shows these variations. At the dataset level, DUC and SummEval show the lowest similarity among human-written summaries across all dimensions. For GUMSum, summaries are more similar to each other. We hypothesize that this is likely due to the constrained annotation guidelines. [^2] https://wiki.gucorpling.org/gum/summarization It is also worth noting that the similarities revealed here are between different human-written summaries for a given instance as opposed to summaries across genres, for which we still expect significant variations, as demonstrated by Liu and Zeldes (2023). Overall, summaries tend to be similar at the syntactic level, less so at the semantic and lexical level. We also observe that LLM outputs show lower diversity (Appendix D), consistently with previous work Giulianelli et al. (2023).
@@ 4.2 Automatic Metrics Fluctuate Substantially at the Instance Level
Given the diversity in human-written summaries, we quantify metric fluctuation at the instance level when using a different set of human-written references. For a metric $M$ and a set of human-written references $R=\{r_{1},r_{2},\dots,r_{N}\}$, we compute $M(r_{i},R-\{r_{i}\})$. Thus, for each document, we score each human-written summary using all the others as the reference set. Figure 4 exemplifies the observed instance-level variability measured by ROUGE-Lavg on the three datasets. For SummEval, we also highlight the scraped original reference in the CNN/DM dataset with a cross. The quality of these references versus the ten crowd-sourced ones is discussed further in Appendix E.
Figure 4: Instance-level variation for ROUGE-Lavg. For every document (shown first 303030, one per line), we plot the score for every human-written reference against all other references (using the same color per source to aid interpretation). The original CNN/DM reference in SummEval is marked by a cross.
@@
For example, the averaged ROUGE-Lavg scores are $28.52_{\pm 5}$, $27.46_{\pm 3}$, $24.88_{\pm 5.3}$ for SummEval, GUMSum, and DUC2004. Given the assumption that human reference summaries are of high quality (i.e., “gold”), metrics should produce high scores. Instead, they do not typically reflect this property.
@@
Understanding the magnitude of such a range might not be obvious. For instance, an increase of 101010 points of BERTScore (typically scoring in the high range of the scale) might indicate a much larger improvement in performance than an increase of 101010 points of ROUGE-1.333The relative dynamics between metrics have been studied by Kocmi et al. (2024) in machine translation. To contextualize the magnitude of variation for each metric, we also report the performance range of summarization systems. Thus, for a model $S$, given its output $o_{i}$ for instance $i$, we score it through $M(o_{i},R)$. Although these values are not directly comparable and should be interpreted with caution due to the use of different reference sets, they help contextualize the magnitude of the results and its potential impact on evaluation. For example, ROUGE-1max assigned to human-written references vary by about 353535 points on average (the green dashed line in Figure 5), while the mean range is less than 202020 points across all model outputs (orange line), and much lower for LLM outputs (blue line). Similarly, LLM summaries exhibit much less variability than human references on all metrics. These findings highlight the significance of variability and suggest that the ranking of summarization models is highly sensitive to the reference set.
@@ 4.3 System Ranking Depends on the Reference(s) for N-gram-based Metrics
We sample $k$ humans-written references ($k\in[1,N]$, where $N$ is the number of references for each document) from all available references for each instance. We then score the outputs of each summarization system using the same set of references. Given $M$ systems $S_{a}$, $S_{b}$, …, $S_{M}$, the metric induces a ranking $S_{a}\succ S_{b}\succ\dots\succ S_{M}$. This process is repeated 100100100 times, yielding 100100100 rankings. We compute the pairwise Kendall rank correlation coefficient Kendall (1938) between such ranks. High correlation indicates that models are similarly ordered, even when different sets of references are used. Figure 6 reports the average correlation for pairs of ranks for each dataset and metric, from using $k$ human-written summaries as references. ROUGEmax is shown in Figure 6, and ROUGEavg is reported in Figure 12 in Appendix G.
@@
In contrast, more semantically-oriented metrics show greater stability. For SummEval, BLEURT shows the highest correlation between ranks, followed by METEOR and BERTScore. BLEURT and METEOR confirm their stability on GUMSum when ranking the LLM outputs. Other metrics (including BERTScore) show low or no correlation on GUMSum, with the exception of ROUGE-1. In all cases, metrics show much higher stability on DUC, for which all average correlations are above 0.70.70.7. We speculate that high stability might stem from an artifact introduced by the short summary length required by the guidelines.
@@
However, the stability varies by metric. ROUGE (especially ROUGEmax) and BLEU tend to have low correlation between ranks. As an example, the ROUGEmax scores require 555-101010 references to reach a level of stability that is comparable to that of BERTScore on SummEval with a single reference. ROUGEavg has a better stability than ROUGEmax, especially with a larger set of references. For example, on SummEval, ROUGE-Lavg has higher stability than BERTScore for $k>3$, while on GUMSum, ROUGE-2avg is the second most stable metric for $k>3$. On all datasets, BLEURT and METEOR remain stable even with a single reference, with METEOR showing stability despite its simplicity.
In general, trends on SummEval are clearer and simpler to interpret than the other two datasets. We speculate that this is due to the larger number of models used (161616 pre-LLM models+444 LLMs on SummEval vs 444 LLMs on GUMSum and DUC). BLEURT, METEOR, and BERTScore show the highest stability, while n-gram-based metrics show low to average correlation between ranks even when multiple references are used. The cases of GUMSum and DUC2004 are more complex to interpret and might be less meaningful given fewer model outputs (i.e., only four LLM outputs). For GUMSum, BLEURT continues to show high inter-rank correlation, with METEOR being the second most stable. BERTScore, on the other hand, shows poor stability. Similar to the case with $k=1$, on DUC2004, all metrics show high stability, likely due to summaries being very short, as dictated by the guidelines.

### v3

@@ 1 Introduction
The assumption behind the use of reference-based metrics is that system outputs that are more similar to the reference(s) are better, due to their “human-likeness” Gehrmann et al. (2023). However, the significant variation in human-written summaries implies that evaluating system outputs against a single or limited set of references has inherent drawbacks. Previous research has extensively looked at correlations between metrics and human judgments in summarization Forde et al. (2024); Mondshine et al. (2025), further exploring the use of multiple references to improve such correlations Lin (2004); Belz and Reiter (2006); Fabbri et al. (2021); Tang et al. (2024) as well as the interpretability and efficiency aspects of such automatic metrics Liu et al. (2023b). However, a much less studied question is the extent to which automatic metrics are sensitive to the choice of human-written reference summaries, as shown in Figure 1. In other words, are these metrics stable across different plausible gold-standard references? If metric scores vary significantly with the selected reference(s), this variation calls into question the reliability of many evaluation practices in the field.
@@ 2 Related Work
Recent advances in Natural Language Generation (NLG) have significantly enhanced the development of summarization systems. However, their evaluation remains an open problem Celikyilmaz et al. (2021); Goyal et al. (2023). Summarization evaluation metrics are broadly categorized into reference-based and reference-free Lloret et al. (2018). Reference-based metrics compare system outputs to human-written reference summaries, relying on methods such as n-gram overlap Lin (2004); Papineni et al. (2002), embedding similarity Ng and Abrecht (2015); Zhao et al. (2019); Zhang et al. (2020), or model-based techniques Peyrard et al. (2017); Scialom et al. (2019); Yuan et al. (2021). In contrast, reference-free summarization metrics do not assume a gold standard Yuan et al. (2021); Vasilyev et al. (2020); Gao et al. (2020); Gigant et al. (2024). More recently, growing research leverages LLMs as evaluators, with or without references Song et al. (2024); Li et al. (2024). In the LLM-as-judge paradigm, evaluations are typically based either on prompting the model to provide judgments Liu et al. (2023a); Bavaresco et al. (2025), or on using its generative probabilities directly Fu et al. (2024).
@@
Meta-evaluation of summarization metrics typically focuses on the extent to which they can be used as a proxy for human evaluation. Reiter and Belz (2009) examined the validity of automatic scores for NLG tasks, while Rankel et al. (2013) focused on ROUGE and its correlation with humans. Peyrard (2019) showed that metrics with reasonable correlation on lower-quality outputs tend to diverge when output quality increases. Caglayan et al. (2020) demonstrated the idiosyncrasies of automatic evaluation metrics, noting that high correlation with human judgments is not sufficient to characterize their reliability. Fabbri et al. (2021) performed a large-scale meta-evaluation of summarization metrics, and found that most metrics have low correlation with human judgments on coherence, while relevance is weakly or moderately correlated. Mondshine et al. (2025) performed a meta-analysis of reference-based, reference-free, and LLM-based summarization metrics focusing on eight languages from four typological families, showing low correlation. They also observed that off-the-shelf LLMs as judges still lag behind other metrics.
While most existing research focused on correlation to human scores, Tang et al. (2024) addressed the challenge of evaluation when a limited number of references is available. They proposed leveraging LLMs to diversify the references, expanding the evaluation coverage and improving the correlation with humans. Their results show that increasing the number of references significantly enhances the reliability of existing evaluation metrics in terms of correlation. However, since LLM outputs tend to show less variability and follow distinct patterns compared to human-produced content Giulianelli et al. (2023); Guo et al. (2024); Shur-Ofry et al. (2024); Reinhart et al. (2025), relying on them to replace human references might introduce biases. Beyond summarization, evaluation of reference variability has also been conducted in tasks such as machine translation (Castilho, 2020; Popović, 2021; Wu et al., 2025) and image captioning (Yi et al., 2020).
@@ 3 Experimental Setup
To quantify the impact of human-written references on the scores of automatic metrics, we leverage multiple elements. For datasets, we use SummEval Fabbri et al. (2021), GUMSum Liu and Zeldes (2023), and DUC2004 Dang and Croft (2004), which contain multiple human-written summaries (§3.1), to assess how different reference summaries affect metric performances. Next, to assess summarization models, we use the existing outputs provided by Fabbri et al. (2021) for SummEval. As these outputs predate LLMs, we additionally collect outputs using LLMs (§3.2) for all three datasets. Lastly, to compute the correlations with humans, we use the human judgments available in SummEval and gather new human ratings for GUMSum on both human and LLM-generated summaries (§3.3). We prioritize GUMSum over DUC2004, as it includes multiple genres beyond news data. Our metric selection is outlined in §3.4. Details on data licensing and codebase are provided in Appendix A.
@@ 3.1 Human-written Summaries
SummEval Fabbri et al. (2021) is built on top of CNN/DM Hermann et al. (2015); Nallapati et al. (2016), containing news articles and human-written highlights. The original authors selected $100$ instances from the test set and supplemented the existing highlights with ten additional reference summaries per instance, obtained via crowd-sourcing Kryscinski et al. (2019).
GUMSum Liu and Zeldes (2023) contains summaries created following general and genre-specific guidelines [^1] https://wiki.gucorpling.org/gum/summarization to function as a substitute for the source Nenkova and McKeown (2011). We focus on the $48$ documents in the dev and test sets, which contain five human-written summaries each Lin and Zeldes (2025) across $12$ genres.
DUC2004 Task1 Dang and Croft (2004) consists of $489$ news documents, most with four references. The guidelines allow the summaries to be in the form of short sentences or lists of keywords. [^2] https://duc.nist.gov/duc2004/tasks.html DUC2004 references are thus extremely concise (only up to $75$ characters). The dataset has played a significant role in summarization research, being part of the annual TREC conference evaluation.
Table 1 provides an overview of the three datasets. We treat all human references in the three datasets as “gold” since they were either authored by experts or validated through review.
@@ 3.2 Model Outputs
Fabbri et al. (2021) collected model outputs for SummEval from $24$ extractive and abstractive summarization systems, which were SOTA between 2017 and 2019. We focus on the $16$ models for which they provided human judgments.
For all datasets, we also include summaries generated by contemporary LLMs. This is crucial given that prior studies demonstrated that evaluation metrics often show lower correlation with high-quality outputs Peyrard (2019); Alva-Manchego et al. (2021). Below, we report a similar pattern for LLMs (§4.4). For consistency purposes, we follow Lin and Zeldes (2025) and use Llama3-3B-Instruct Hermann et al. (2015), Qwen-2.5-7B-Instruct Yang et al. (2025), Claude-3.5 Anthropic (2024), and GPT-4o OpenAI (2024). For each LLM, we generate a single summary. We emphasize LLM variety over multiple generations. Details on the generation parameters and prompts are reported in Appendix A.
@@ 3.3 Human Judgments
SummEval (Fabbri et al., 2021) contains expert judgments that assess summaries based on four criteria: coherence, consistency, fluency, and relevance, using a Likert scale of 11-55 Likert (1932).
To measure how well automatic metrics align with human judgments beyond the news domain, and to study whether findings on pre-LLM models align with those on LLM outputs, we collect a new set of human judgments using the same criteria on the $48$ GUMSum documents. We hired three Master’s students in Computational Linguistics and tasked them to evaluate four LLM outputs (§3.2) and five human references (§3.1), following Fabbri et al. (2021)’s criteria.
[…]
claude $\mathbf{4.75_{0.45}}$ $4.48_{0.65}$ $\mathbf{4.82_{0.42}}$ $\mathbf{4.26_{0.71}}$ $\mathbf{0.17_{0.38}}$ $\mathbf{0.03_{0.18}}$
gpt4o $4.42_{0.61}$ $\mathbf{4.61_{0.63}}$ $4.74_{0.45}$ $4.07_{0.65}$ $0.11_{0.32}$ $0.12_{0.32}$
Qwen $4.66_{0.56}$ $4.33_{0.84}$ $4.68_{0.52}$ $4.23_{0.75}$ $0.16_{0.37}$ $0.12_{0.33}$
Llama3 $4.71_{0.51}$ $4.14_{0.97}$ $4.78_{0.42}$ $4.09_{0.86}$ $0.12_{0.32}$ $0.20_{0.40}$
humans $54_{0.57} 4.4$$48_{0.71}$ $4.70_{0.51}$ $4.22_{0.69}$ $0.09_{0.28}$ $0.10_{0.31}$
Table 2: Human judgments on GUMSum: LLM vs. human-written summaries. Above-human performances are highlighted in blue.
LLM-generated and human-written summaries were anonymized and shuffled. We also asked the evaluators to pick one best and one worst summary for each document.
Table 2 reports the results. Claude scored the best overall. GPT-4o gets the highest consistency but the lowest coherence and relevance, and is thus the least picked LLM. Interestingly, LLM outputs typically receive higher scores than human-written references. In line with previous work, e.g., Zhang et al. (2024), this finding has significant implications for reference-based evaluation and calls into question the use of potentially lower-quality references for assessing high-quality outputs Noh et al. (2024).
@@ 3.4 Evaluation Metrics
We consider several reference-based metrics, chosen to balance popularity and diversity. All metrics range in 0100100. Appendix B provides details.
@@
BLEU Papineni et al. (2002) is an n-gram metric primarily used for translation. It is precision-based and incorporates a brevity penalty. With multiple references, the n-gram count is clipped at the maximum count of n-grams in a single reference, and the length of the reference closest in size to the hypothesis is considered.
@@
To gain insights into recent metric usage, Figure 2 summarizes the percentage of summarization papers from recent ACL, EMNLP, and INLG proceedings (detailed in Appendix C). We found that reference-based metrics are still the most popular, with ROUGE in 7979% of papers, followed by BERTScore (4444%). Our preliminary search also shows that LLM-as-a-judge evaluators are mainly (i.e., in 6666% of the cases) used without references, placing them outside the scope of this paper.
@@ 4.1 Human-written Summaries are Diverse
Figure 3 shows these variations. At the dataset level, DUC and SummEval show the lowest similarity among human-written summaries across all dimensions. For GUMSum, summaries are more similar to each other. We hypothesize that this is likely due to the constrained annotation guidelines. It is also worth noting that the similarities revealed here are between different human-written summaries for a given instance as opposed to summaries across genres, for which we still expect significant variations, as demonstrated by Liu and Zeldes (2023). Overall, summaries tend to be similar at the syntactic level, less so at the semantic and lexical level. We also observe that LLM outputs show lower diversity (Appendix D), consistently with previous work Giulianelli et al. (2023).
@@ 4.2 Automatic Metrics Fluctuate Substantially at the Instance Level
Given the diversity in human-written summaries, we quantify metric fluctuation at the instance level when using a different set of human-written references. For a metric $M$ and a set of human-written references $R=\{r_{1},r_{2},\dots,r_{N}\}$, we compute $M(r_{i},R-\{r_{i}\})$. Thus, for each document, we score each human-written summary using all the others as the reference set.
Figure 5: Instance-level variation for ROUGE-Lavg. For every document (shown first $30$, one per line), we plot the score for every human-written reference against all other references (using the same color per source to aid interpretation). The original CNN/DM reference in SummEval is marked by a cross.
Figure 5 exemplifies the observed instance-level variability measured by ROUGE-Lavg on the three datasets. For SummEval, we also mark the original reference (scraped highlights, §3.1) in the CNN/DM dataset with a cross. The quality of these scraped references versus the ten later crowd-sourced ones is discussed further in Appendix E.
@@
For example, the averaged ROUGE-Lavg scores are $28.52_{\pm 5}$, $27.46_{\pm 3}$, $24.88_{\pm 5.3}$ for SummEval, GUMSum, and DUC2004. Given the assumption that human reference summaries are of high quality, metrics should produce high scores. Instead, they do not typically reflect this property.
@@
Understanding the magnitude of such a range might not be obvious. For instance, an increase of $10$ points of BERTScore (typically scoring in the high range of the scale) might indicate a much larger improvement in performance than an increase of $10$ points of ROUGE-1.333The relative dynamics between metrics have been studied by Kocmi et al. (2024) in machine translation. To contextualize the magnitude of variation for each metric, we also report the performance range of summarization systems. Thus, for a model $S$, given its output $o_{i}$ for instance $i$, we score it through $M(o_{i},R)$. Although these values are not directly comparable and should be interpreted with caution due to the use of different reference sets, they help contextualize the magnitude of the results and its potential impact on evaluation. For example, ROUGE-1max assigned to human-written references varies by about $35$ points on average (the green dashed line in Figure 4), while the mean range is less than $20$ points across all model outputs (orange line), and much lower for LLM outputs (blue line). Similarly, LLM summaries exhibit much less variability than human references on all metrics. These findings highlight the significance of variability and suggest that the ranking of summarization models is highly sensitive to the reference set.
@@ 4.3 System Ranking Depends on the Reference(s) for N-Gram-Based Metrics
We sample $k$ human-written references ($k\in[1,N]$, where $N$ is the number of references for each document) from all available references for each instance. We then score the outputs of each summarization system using the same set of references. Given $M$ systems $S_{a}$, $S_{b}$, …, $S_{M}$, the metric induces a ranking $S_{a}\succ S_{b}\succ\dots\succ S_{M}$. This process is repeated $100$ times, yielding $100$ rankings. We compute the pairwise Kendall rank correlation coefficient Kendall (1938) between such ranks. High correlation indicates that models are similarly ordered, even when different sets of references are used. [^4] Note that, for $k>1$, references in different reference sets might overlap, artificially increasing the observed correlation between rankings. Figure 6 reports the average correlation for pairs of ranks for each dataset and metric, from using $k$ human-written summaries as references. ROUGEmax is shown in Figure 6, and ROUGEavg is reported in Figure 12 in Appendix G.
@@
In contrast, more semantically-oriented metrics show greater stability. For SummEval, BLEURT shows the highest correlation between ranks, followed by METEOR and BERTScore. BLEURT and METEOR confirm their stability on GUMSum when ranking the LLM outputs. Other metrics (including BERTScore) show low or no correlation on GUMSum, with the exception of ROUGE-1. In all cases, metrics show much higher stability on DUC, for which all average correlations are above $0.7$. We speculate that high stability might stem from an artifact introduced by the short summary length required by the guidelines.
@@
However, the stability varies by metric. ROUGE (especially ROUGEmax) and BLEU tend to have low correlation between ranks. As an example, the ROUGEmax scores require 55-$10$ references to reach a level of stability that is comparable to that of BERTScore on SummEval with a single reference. ROUGEavg has a better stability than ROUGEmax, especially with a larger set of references. For example, on SummEval, ROUGE-Lavg has higher stability than BERTScore for $k>3$, while on GUMSum, ROUGE-2avg is the second most stable metric for $k>3$. On all datasets, BLEURT and METEOR remain stable even with a single reference, with METEOR showing stability despite its simplicity.
In general, trends on SummEval are clearer and simpler to interpret than the other two datasets. We speculate that this is due to the larger number of models used ($16$ pre-LLM models+$4$ LLMs on SummEval vs $4$ LLMs on GUMSum and DUC). BLEURT, METEOR, and BERTScore show the highest stability, while n-gram-based metrics show low to average correlation between ranks even when multiple references are used. The cases of GUMSum and DUC2004 are more complex to interpret and might be less meaningful given fewer model outputs (i.e., only four LLM outputs, which might increase the observed noise). For GUMSum, BLEURT continues to show high inter-rank correlation, with METEOR being the second most stable. BERTScore, on the other hand, shows poor stability. Similar to the case with $k=1$, on DUC2004, all metrics show high stability, likely due to summaries being very short, as dictated by the guidelines.